The length of the base of a rectangular prism is twice its width. The perimeter of the base is 42 yards. If the volume of the prism is 980 cubic yards, find the height of the prism.


1
Expert's answer
2022-03-21T16:03:50-0400

The volume of a rectangular prism is the product of its three dimensions V=l*w*h, where l is length, w is width and h is height. It means that "h=\frac{V}{lw}" . For the first, lets find length and width of the prism. The perimeter P=2(l+w) and in our case l=2w, so 42=2(2w+w)=6w, from what w=7 yards and l=14 yards. Finally, "h=\frac{V}{lw}=\frac{980}{14*7}=\frac{980}{98}=10" yards.

Answer: h=10 yards.